OpenStudy (anonymous):

please consumer math help you owe 2,264.96 on a credit card at a 14.7% APR you pay 500.00/month. How many more months does it take to pay off your balance?

OpenStudy (wolf1728):

The formula for determining months remaining is attached. (Yes it is THAT complicated). http://1728.org/loanfrm3.htm annual rate = 14.7% monthly rate = 1.225% = .01225 Months = [log(1+(rate/(pmt/princ)-rate))] / [log(1+rate)] Months = [log(1+(.01225/(500/2264.96)-.01225)] / [log(1.01225)] Months = [log(1+(0.05549152-.01225)] / [log(1.01225)] Months = [log(1+ (0.04324152) / 0.0052877854] Months = 0.0183848632 / 0.0052877854 Months = 3.4768549968 This is NOT (I mean NOT an easy formula to work with!!) The above calculations are the second time I have calculated this and it is STILL wrong. There is a calculator for this located here: http://1728.org/calcloan.htm which says the time is 4.692 MONTHS